A company currently pays a dividend of $3.5 per share (D0 = $3.5). It is estimated that the company's dividend will grow at a rate of 20% per year for the next 2 years, and then at a constant rate of 6% thereafter. The company's stock has a beta of 1.5, the risk-free rate is 3.5%, and the market risk premium is 6%. What is your estimate of the stock's current price? Round your answer to the nearest cent.
Answer #1

Last Dividend, D0 = $3.50

Growth rate for next 2 years is 20% and a constant growth rate (g) of 6% thereafter

D1 = $3.50 * 1.20 = $4.20
D2 = $4.20 * 1.20 = $5.04
D3 = $5.04 * 1.06 = $5.3424

Required Return, rs = Risk-free Rate + Beta * Market Risk Premium
Required Return, rs = 3.50% + 1.50 * 6.00%
Required Return, rs = 12.50%

Stock Price in 2 years, P2 = D3 / (rs - g)
Stock Price in 2 years, P2 = $5.3424 / (0.1250 - 0.0600)
Stock Price in 2 years, P2 = $5.3424 / 0.0650
Stock Price in 2 years, P2 = $82.1908

Current Stock Price, P0 = $4.20/1.1250 + $5.04/1.1250^2 + $82.1908/1.1250^2
Current Stock Price, P0 = $72.66